OpenStudy (anonymous):

Determine the radius and interval of convergence of the power series sum from n=1 to infinity of (x-2)^n/n(4)^n

OpenStudy (anonymous):

ratio test?

OpenStudy (anonymous):

can you show me how to do it please. its giving me a hard time

OpenStudy (anonymous):

with some practice you can probably see from your eyeballs that it will converge if \(|x-2|<4\)

OpenStudy (anonymous):

can you please show me in full steps and I will follow your note and practice it?

OpenStudy (anonymous):

ok we are going to put \[a_n=\frac{(x-2)^n}{n\times 4^n}\] and then compute \[\frac{a_{n+1}}{a_n}\]

OpenStudy (anonymous):

of course you don't write a compound fraction, you write \[\frac{(x-2)^{n+1}}{(n+1)\times 4^{n+1}}\times \frac{n\times 4^n}{(x-2)^n}\]

OpenStudy (anonymous):

then a whole raft of cancellation gives \[\frac{n(x-2)}{4(n+1)}\]

OpenStudy (anonymous):

in fact you can really just ignore the \(x-2\) for a moment. now we take \[\lim_{n\to \infty}\frac{n}{4(n+1)}\] which is pretty clearly \(\frac{1}{4}\)

OpenStudy (anonymous):

at least i think it is clear, if not, let me know

OpenStudy (anonymous):

this will therefore converge if \(|x-2|<4\) because if the limit is less than one, you can compare it to a geometric series and see that it converges you don't have to worry about that part, that is part of the proof that the ratio test works

OpenStudy (anonymous):

in other words, if \[\lim_{n\to \infty}\frac{a_{n+1}}{a_n}=L\] then the series converges is \(|x-a|<\frac{1}{L}\)

OpenStudy (anonymous):

damn i did not mean \(x-a\) that used \(a\) twice i guess i should have used \(|x-b|\) or something

OpenStudy (anonymous):

n(x−2)/4(n+1) but isn't it going to be |R|<4 which leads to |x-2|<4 and this become |x|<6 and radius become 6 ?

OpenStudy (anonymous):

in any case your entire job is to compute \[\lim_{n\to \infty}\frac{a_{n+1}}{a_n}\] and then the radius of convergence will be the reciprocal of your answer

OpenStudy (anonymous):

oh bad algebra mistake there !

OpenStudy (anonymous):

\(|x-2|<4\) does not mean \(|x|<6\)

OpenStudy (anonymous):

\[|x-2|<4\iff -4<x-2<4\iff -2<x<6\]

OpenStudy (anonymous):

and this the radius of convergence right? what about the interval ?

OpenStudy (anonymous):

the word "radius" will be clearer when you work with complex variables

OpenStudy (anonymous):

ok let me say it correctly the radius is 4, the "interval" is actually the interval, the solution to \(|x-2|<4\) which is \[-2<x<6\]

OpenStudy (anonymous):

you also might be asked to check the endpoints of the interval as well

OpenStudy (anonymous):

because in this case the interval of convergence is actualy \[[-2,6)\] that is, it converges if \(x=-2\) but does not converge if \(x=6\)

OpenStudy (anonymous):

do you know how to check that?

OpenStudy (anonymous):

I'm not quite sure. Im taking calc for the first time and its a pain in the butt when your taking it at the summer too many materials and too little time

OpenStudy (anonymous):

yeah summer courses like a forced march through math, not often a good idea

OpenStudy (anonymous):

but in this case it is not too bad, replace \(x\) by \(6\) and see what you get

OpenStudy (anonymous):

you can almost do it in your head \[\sum\frac{4^n}{n\times 4^n}=\sum\frac{1}{n}\] which is well known to diverge

OpenStudy (anonymous):

so the 6 is out

OpenStudy (anonymous):

then replace \(x\) by \(-2\) and get \[\sum\frac{(-4)^n}{n\times 4^n}\] which is an alternating series, and since the terms go to zero, it does converge

OpenStudy (anonymous):

that is why the exact answer for the interval of convergence is \[[-2,6)\]

OpenStudy (anonymous):

if you have any questions let me know

OpenStudy (anonymous):

so if this question comes in my exam next week. and I did every exct thing, I would get 100 out 100 in this question

OpenStudy (anonymous):

yes but how likely is that?

OpenStudy (anonymous):

we can do another if you like

OpenStudy (anonymous):

I mean if i get the same problem and I copied this answer above, do you think I would get a full mark based on the answers above?

OpenStudy (anonymous):

Thanks you very much that was helpful.

OpenStudy (anonymous):

yes, i think so especially if you checked the endpoints

OpenStudy (anonymous):

you are the best :)

OpenStudy (anonymous):

why thank you hope it was more or less clear

OpenStudy (anonymous):

I acully have another question if you doint mind

OpenStudy (anonymous):

let R be the region enclosed by the curves y=18-x^2 and y=x^2. use the following method to set up an integral, but don't evaluate? and the volum of the solid obtained by rotating R about the x-axis ?